Write the slope-intercept equation for the line that passes through (14, -2) and is perpendicular to 7x – 10y = 18
Answer by alicealc(293)

7x - 10y = 18
-10y = -7x + 18
divide all by -10
y = 7/10 x - 18/10
the slope of this equation is 7/10

if 2 lines are perpendicular then
slope 1 * slope 2 = -1
7/10 * slope 2 = -1
slope 2 = -1 * 10/7 = -10/7

line equation for the line with slope m and passes through (x1, y1):
y - y1 = m*(x - x1)

m = -10/7
x1 = 14
y1 = -2
y - (-2) = -10/7 * (x - 14)
y + 2 = -10/7 x + 140/7
multiply all by 7:
7y + 14 = -10x + 140
10x + 7y + 14 - 140 = 0
so the equation will be:
10x + 7y - 126 = 0